Given that, for AgI, at 25°C, Ksp = 1.5×10-16:
From a 1.0×10-6-M NaI solution, precipitation of AgI just begins at what concentration of added Ag+?
Answer #1

Solubility equilibrium of AgI is

AgI(s) <--------> Ag+(aq) + I-(aq)

Ksp = [Ag+][I-] = 1.5 ×10-16

substituting given concentration of I-

[Ag+][I-] = 1.5 ×10-16 M2

[Ag+] × 1.0 ×10-6M = 1.5 × 10-16M2

[Ag+] = 1.5 × 10-10M

Therefore,

Concentration of Ag+ required to start precipitation of AgI = 1.5 × 10-10M